• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name D. Lambert

    Log entry time 09:08:13 on June25,2004

    Entry number 126812

    keyword=IOC reboot

    rebooted IOC about 8:50.  white alarm was not present, strip chart readings just flattened, with temperature reading on TV monitor climbing.  It seems that things went white while beam was on at a low current, but off during the frozen time.  Paged Kathy, went to manual heater to stabilize during reboot, everything returned to normal.  alarm settings for loops 1 and 3 reverted to old settings, and were set to no alarm.  Don't know why white alarm was not present.